# Revisiting the Predictability of Performative, Social Events

### v1

@@ 4 Structural Results
In the previous section, we illustrated how one can find predictors $f$ that actively shape the data $(x,y)\sim{\cal D}(f)$, yet still make predictions that are computationally indistinguishable from the true outcomes. Here, we take a step further and ask: are these computationally indistinguishable predictors truly valuable? To answer this question, we analyze the relationship between the previous historical desiderata (Definition 1.1) and the core solution concepts in performative prediction. We review these below.
@@
$\operatorname*{\mathbb{E}}_{(x,y)\sim{\cal D}(f_{\mathrm{po}}),p\sim f_{\mathrm{po}}(x)}\ell(p,y)\leqslant\min_{h\in{\cal H}}\operatorname*{\mathbb{E}}_{(x,y)\sim{\cal D}(h)}\ell(h(x),y).$
@@
The following theorem is the main result of this section, showing how performatively multicalibrated predictors as per Definition 1.1 are also performatively stable with respect to standard losses like squared error. To simplify our presentation, we make the further assumption that outcomes $y$ are binary. [^8] Given recent advances [Gopalan et al., 2024, Lu et al., 2025], it is not hard to generalize this loss minimization guarantee to the case where $y$ is real-valued or to other loss functions. The main idea is to produce outcome indistinguishable predictions of a vector consisting of multiple statistics or moments of the outcome $y$. These extensions are interesting, but also somewhat involved and not essential to the main conceptual point of this section relating indistinguishability and stability. We omit them to keep our presentation short.
@@
If $\ell$ is the squared loss $\ell(p,y)=\frac{1}{2}(1-p)^{2}$, then
@@
These results strengthen our previous understanding of performative stability. In the initial work on performative prediction, stable predictors were only known to be efficiently computable under strong regularity conditions on the distribution map ${\cal D}(\cdot)$ [Perdomo et al., 2020, Mendler-Dünner et al., 2020, Drusvyatskiy and Xiao, 2023]. Specifically, ${\cal D}_{y}(x,p)$ is assumed to be sufficiently Lipschitz in the forecast $p$. This restriction rules out common settings where decision makers choose actions that influence outcomes if the forecast $p$ is above or below some threshold. In education, for instance, counselors at a school often assign extra attention to students by examining whether their predicted probabilities are below some fixed threshold [Perdomo et al., 2023]. This thresholding implies that the distribution map is not Lipschitz.
@@ 5 Suboptimality of Perfectly Calibrated Performative Predictions
    For any function, $h:\mathcal{X}\rightarrow[0,1]$, $\operatorname*{\mathbb{E}}_{(x,p,y)\sim{\cal D}(f)}(p-y)^{2}\geqslant\operatorname*{\mathbb{E}}_{(x,y)\sim{\cal D}(h)}(y-h(x))^{2}-{\cal O}(\varepsilon).$ Hence,
@@
Figure 1: A visualization of the distribution map for the construction in Theorem 5.1. The solid black lines describe the probability that $y=1$ given that the prediction is equal to $p$. The blue lines indicate the probability that $y=1$ if one deploys the randomized predictor $f_{r}$ mixes between $1/2$ with probability $\lambda$ and $1/2+\varepsilon$ with probability $1-\lambda$ for the whole range of $\lambda\in[0,1]$. Fixed points, $\operatorname*{\mathbb{E}}_{{\cal D}(f)}[y]=\operatorname*{\mathbb{E}}_{{\cal D}(f)}[f(x)]$ are those that cross the dotted diagonal line.
@@
since $y$ is either 1 or 0 and $p$ is always nearly 1/2. The performative risk for this problem is a quadratic function in $p$ that is maximized at $p=1/2$. Yet the performatively optimal solutions are to predict either 0 or 1. For these choices, we would have the best possible performative risk of .01.01.01.01. ∎
@@ Acknowledgements
We would like to thank Aaron Roth and Tijana Zrnic for helpful comments and discussion.
@@ Appendix A The K29 Algorithm
To apply the algorithm to specific function classes ${\cal H}$, the only remaining todo is to show how to construct kernels $k$ with corresponding RKHS ${\cal F}$ such that ${\cal H}\subseteq{\cal F}$ and to show that both kernel evaluations $k((x,p),(x,p))$ and function norms $\|h\|_{{\cal F}}$ are uniformly bounded. This yields the ${\cal O}(\sqrt{T})$ regret bound. See Dwork et al. [2024] for examples on how to do this for common function classes.

### v2

@@ 4 Structural Results
In the previous section, we illustrated how one can find predictors $f$ that actively shape the data $(x,y)\sim{\cal D}(f)$ and simultaneously make predictions that are computationally indistinguishable from the true outcomes. Here, we take a step further and ask: are these computationally indistinguishable predictors useful (in a risk minimization sense)? What is the relationship between calibration and loss minimization in this outcome performative context?
To answer these, we analyze the relationship between the previous historical desiderata (Definition 1.1) and the core solutions in performative prediction. We review these below.
@@
$\operatorname*{\mathbb{E}}_{(x,p,y)\sim{\cal D}(f_{\mathrm{po}})}\ell(p,y)\leqslant\min_{h\in{\cal H}}\operatorname*{\mathbb{E}}_{(x,y)\sim{\cal D}(h)}\ell(h(x),y).$
@@
The following theorem is the main result of this section, showing how performatively multicalibrated predictors as per Definition 1.1 are also performatively stable with respect to standard losses like squared error. To simplify our presentation, we make the further assumption that outcomes $y$ are binary. [^6] Given recent advances [Gopalan et al., 2024, Lu et al., 2025], one can generalize this loss minimization guarantee to the case where $y$ is real-valued or to other loss functions. The main idea is to produce outcome indistinguishable predictions of a vector consisting of multiple statistics or moments of the outcome $y$. These extensions are interesting, but also somewhat involved and not essential to the main conceptual point of this section relating indistinguishability and stability. We omit them to keep our presentation short.
@@
If $\ell$ is the squared loss $\ell(p,y)=\frac{1}{2}(y-p)^{2}$, then
@@
These results strengthen our previous understanding of performative stability. In the initial work on performative prediction, stable predictors were only known to be efficiently computable under strong regularity conditions on the distribution map ${\cal D}(\cdot)$ [Perdomo et al., 2020, Mendler-Dünner et al., 2020, Drusvyatskiy and Xiao, 2023] and the loss function $\ell$. Specifically, ${\cal D}_{y}(x,p)$ is assumed to be sufficiently Lipschitz in the forecast $p$ and $\ell$ is pointwise strongly convex and smooth. This restriction on ${\cal D}(\cdot)$ rules out common settings where decision makers choose actions that influence outcomes if the forecast $p$ is above or below some threshold. In education, for instance, counselors at a school often assign extra attention to students by examining whether their predicted probabilities are below some fixed threshold [Perdomo et al., 2025]. This thresholding implies that the distribution map is not Lipschitz.
@@ 5 Suboptimality of Perfectly Calibrated Performative Predictions
    Furthermore, for any function, $h:\mathcal{X}\rightarrow[0,1]$,
    $\operatorname*{\mathbb{E}}_{(x,p,y)\sim{\cal D}(f)}(p-y)^{2}\geqslant\operatorname*{\mathbb{E}}_{(x,y)\sim{\cal D}(h)}(y-h(x))^{2}-{\cal O}(\varepsilon).$
    Hence,
@@
Figure 1: A visualization of the distribution map for the construction in Theorem 5.1. The solid black lines describe the probability that $y=1$ given that the prediction is equal to $p$. The blue dotted line indicates the probability that $y=1$ if one deploys the randomized predictor $f_{r}$ that mixes between $1/2$ with probability $\lambda$ and $1/2+\varepsilon$ with probability $1-\lambda$ for the whole range of $\lambda\in[0,1]$. Fixed points, $\operatorname*{\mathbb{E}}_{{\cal D}(f)}[y]=\operatorname*{\mathbb{E}}_{{\cal D}(f)}[f(x)]$ are those that cross the dotted diagonal line. Because $\mathrm{Pr}[y=1\mid f(x)=p]$ is discontinuous in $p$, there is no deterministic prediction $p$ that is calibrated, i.e $p=\mathrm{Pr}[y=1\mid f(x)=p]$. The only calibrated predictors are those that randomize between $1/2$ and $1/2+\varepsilon$. While calibrated, these maximize squared error since they induce $y$ to be a fair coin toss.
@@
since $y$ is either 1 or 0 and $p$ is always nearly 1/2. The performative risk for this problem is a quadratic function in $p$ that is maximized at $p=1/2$. Yet the performatively optimal solutions predict either 0 or 1. These achieve the best possible performative risk of .01.01.01.01. ∎
@@
6 Discussion and Future Work
Prediction algorithms are now commonplace in important social domains, including healthcare, education, and public administration. Performativity abounds. By shaping our decision-making processes, these prediction algorithms influence the outcomes we see. Drawing on a long intellectual history, our work revisits core methodological questions regarding the design and evaluation of predictors in these domains.
On the design side, we introduced new algorithms that can overcome social feedback and produce rigorously calibrated predictions of social events. These results resolve algorithmic questions left open by Simon [1954] and Grunberg and Modigliani [1954]. On the evaluation side, our contributions are mostly conceptual. We illustrate how traditional desiderata imported from supervised learning— calibration and its different variants — mean something quite different in these performative contexts where predictions shape outcomes. These insights call into question claims regarding the utility and significance of public forecasts of social events that stake their validity on calibration [Silver, 2019].
There are a number of interesting directions for future work in this area. For one, we prove our results for the simplest case of performative prediction: (stateless) outcome performativity. It would be interesting to consider whether these results carry over to richer, stateful domains where the outcomes we see don’t just depend on the predictions we make today, but also on the predictions we’ve made in the past [Brown et al., 2022]. It would also be valuable to fully understand whethere these results carry over to the setting where both feature and outcomes are performative.
@@ Acknowledgements
We would like to thank Aaron Roth, Tijana Zrnic, and the anonymous ICML reviewers for helpful comments and discussion.
@@ Appendix A The K29 Algorithm
To apply the algorithm to specific function classes ${\cal H}$, the only remaining todo is to show how to construct kernels $k$ with corresponding RKHS ${\cal F}$ such that ${\cal H}\subseteq{\cal F}$ and to show that both kernel evaluations $k((x,p),(x,p))$ and function norms $\|h\|_{{\cal F}}$ are uniformly bounded. This yields the ${\cal O}(\sqrt{T})$ regret bound. See Dwork et al. [2025] for examples on how to do this for common function classes and Perdomo and Recht [2025] for a introduction to these techniques.